OpenStudy (anonymous):

Bottled water in a certain vending machine costs $1.50. This price is a markup from the wholesale cost of $0.20. What is the markup as a percent?

OpenStudy (mayankdevnani):

87%

OpenStudy (anonymous):

surely its 750%?

OpenStudy (mayankdevnani):

@tiname

OpenStudy (anonymous):

0.2*750% =1.5

OpenStudy (mayankdevnani):

how you got it??? @Razzputin

OpenStudy (anonymous):

surely it is

OpenStudy (anonymous):

\[\frac{ 1.5 }{ \frac{ 0.2 }{ 100 } } = 750\]

OpenStudy (anonymous):

you take the marked up price and divide it by the origional price over 100 to get a percentage.

OpenStudy (anonymous):

how did you get your answer?

OpenStudy (mayankdevnani):

no, it would be like that Subtract original from markup (1.50-.20) = 1.30 and then divide by markup (1.50) which gets you 87%.

OpenStudy (anonymous):

when you mark something up you do it on the origional item price in this case the 0.2 not on the final marked up price.

OpenStudy (anonymous):

do you disagree?